Two coils have the same number of circular turns and carry the same current. Each rotates in a magnetic field acting perpendicularly to its axis of rotation. Coil 1 has a radius of 4.8 cm and rotates in a 0.15-T field. Coil 2 rotates in a 0.32-T field. Each coil experiences the same maximum torque. What is the radius (in cm) of coil 2?


1
Expert's answer
2021-03-23T08:15:20-0400
"\tau_a=\tau_b\\N
_o
B
_a
I_
o
A_
a
=
N_
o
B_
b
I_
o
A_
b"

"B_
a
A_
a
=
B
_b
A_
b
 \\ \pi B_aR_a^2=\pi B_bR_b^2\\"

"(0.15)(4.8)^2=(0.32)R_b^2\\R_b=3.29\ cm"